OpenStudy (anonymous):

Is it C?? Manny walks for 30 min at a speed of 5 km/h. How far does Manny walk? A. 2.5 km B. 7.5 km C. 75 km D. 150 km Is it C?

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (butterflydreamer):

not quiteee.

OpenStudy (thomas5267):

You can walk 5 km in one hour. What is the distance you can walk in half an hour?

OpenStudy (sissyedgar):

I agree with C because... 30 minutes is half an hour so you would do 5 / 2 which = 2.5. 2.5 * 30 = 75! But I could be wrong

OpenStudy (anonymous):

i did 30x5=250/2=75

OpenStudy (butterflydreamer):

So the question asks for his DISTANCE in 30 minutes! And the speed they give you is 5km/hr (note: 1 hr = 60 minutes)

OpenStudy (butterflydreamer):

so if he did 5 km in 60 minutes and we want ? km in 30 minutes You simply divide by two :)

OpenStudy (anonymous):

2x30

OpenStudy (anonymous):

5-3=2 so its 2 and 30

OpenStudy (anonymous):

it is either c or d though

OpenStudy (butterflydreamer):